OpenStudy (anonymous):

Kyra observed that 85% of the students at the library liked reading mystery stories. If 30 students don't like reading mysteries the number of students at the library who like mysteries were?

OpenStudy (anonymous):

170

OpenStudy (anonymous):

100 - 85 = 15% so 30 people is 15% of the class 85/15 = 5.666 30 * 5.666 =170 people rounded

OpenStudy (anonymous):

what bannemohr said

OpenStudy (imstuck):

or you could figure it as 30 is 15% of what number of people total? That comes out to 200, then you just subtract the 30 who don't like mysteries from the 200 to get the number of people who do like mysteries. Either way...